OpenStudy (anonymous):

given that the rectangle above has an area of 48 square units, find m angle klj. round the answer to the nearest hundreth. a. 18.43 degrees b. 36.87 degrees c. 41.41 degrees d. 48.59 degrees

OpenStudy (anonymous):

@pitamar

OpenStudy (anonymous):

ok, so first of all KJ = LM = 6, right?

OpenStudy (anonymous):

ok yea

OpenStudy (anonymous):

And similarly we can say KL = JM. Let's call them \(x\) So we can say:  Now we know that the area of the rectangle is \(48_{cm^2}\) So \(6_{cm} \cdot x = 48_{cm^2}\) can you find \(x)?

OpenStudy (anonymous):

*can you find \(x\)?

OpenStudy (anonymous):

42?

OpenStudy (anonymous):

no because it's not 6 + x = 48 it is 6*x = 48 you have to divide both sides by 6. we get $$ x = \frac{48_{cm^2}}{6_{cm}} = \frac{48_{cm}}{6} $$ what is 48/6?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

right, so \(x = 8_{cm}\) Now let's redraw this: